OpenStudy (anonymous):

Help please!! Much Appreciated! (: Solve each system by any method of your choice: 2x-6y=16 3x+18y=-30

OpenStudy (dumbcow):

multiply top equation by 3 .... this way you can eliminate "y" 6x -18y = 48 3x +18y = -30 -------------- 9x +0 = 18 --> x = 2 plug x=2 into either equation to solve for y

OpenStudy (anonymous):

So: 3x+18(2)=-30 3x+36=-30 -36 -36 3x=-6 = -2

OpenStudy (dumbcow):

haha not quite just you just happened to get the right answer

OpenStudy (dumbcow):

3(2) +18y = -30 6 +18y = -30 -6 -6 ------------- 18y = -36 y = -2

OpenStudy (anonymous):

haha(: well that's weird :P ...so how would you do it?

OpenStudy (anonymous):

Above method is called elimination method. Here is substitution method: 1) From equation 6x-18y=48 we can deduce x=(48+18y)/6 that will be x=8+3y Now u can substitute the value of x=8+3y in second equation (i.e 3x+18y=-30) 2) On substitution 3(8+3y)+18y=-30 =>24+9y+18y=-30 =>27y=-54 =>y=-2 3)Now by substuting y in anu of the equation u can get the value of x

OpenStudy (primeralph):

cutie less than 3 ? okay.

OpenStudy (anonymous):

Okay, that's how I was doing it at first then got mixed up. Thanks! I really appreciate all of your guy's help!! GBU!!! :) @primeralph Do you mean the answer is less than 3?

OpenStudy (primeralph):

No, your name says cutie less than 3.

OpenStudy (anonymous):

Oh, haha! @primeralph very funny:P